# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v1

@@ 4.2. Extension of kernel EDMD to control systems
In this section, we recap the extension of kEDMD as introduced in Subsection 4.1 to systems with inputs introduced in BoldPhil24 ; BoldScha25 . The kernel-EDMD surrogate model is defined for control-affine discrete-time systems
@@
with locally Lipschitz-continuous maps $g_{0}:\Omega\rightarrow\mathbb{R}^{n}$ and $G:\Omega\rightarrow\mathbb{R}^{n\times m}$ with Lipschitz constants $L_{g_{0}},L_{G}$ on $\Omega$ and control input $u$ restricted to a compact set $\mathbb{U}\subset\mathbb{R}^{m}$ containing the origin in its interior. We denote by $g_{i}:\Omega\rightarrow\mathbb{R}^{n}$ the $i$-th column of the matrix-valued map $G$. Moreover, $g_{0}(0)=0$ is assumed to render the origin an equilibrium of the system (22) with $u=0$. For the relation to continuous-time systems, we refer the interested reader to (BoldPhil24, , Rem. 4.2) and StraScha25 , where it is shown that the uniform and proportional error bounds presented below are preserved for fixed $\varepsilon$ and sufficiently small time step $\Delta t>0$ if only the continuous-time system is control affine.
@@
Assumption 5 (Data requirements).
Let $\mathcal{X}$ defined by (19) with $x_{1}=0$ and a cluster radius $r_{\mathcal{X}}\geq 0$ be given. Then, for each $i\in[1:d]$, let $d_{i}\geq m+(1-\delta_{1i})$ and data triplets $(x_{ij},u_{ij},x_{ij}^{+})$, $j\in[1:d_{i}]$, with pairwise distinct $u_{ij}$ be given such that ${x}_{ij}\in\mathcal{B}_{r_{\mathcal{X}}}(x_{i})\cap\Omega$, $x_{ij}^{+}=F(x_{ij},u_{ij})$, and
$\operatorname{rank}([u_{i1}\mid\ldots\mid u_{id_{i}}])=m.$ (23)
Assumption 5 requires $D:=\sum_{i=1}^{d}d_{i}$ data triplets. We call $x_{i}$ a virtual observation (or cluster) point to emphasize that no samples at $x_{i}$ are required. We define the matrices
$U_{i}:=\begin{bmatrix}1&\dots&1\\ u_{i1}&\dots&u_{id_{i}}\end{bmatrix}\in\mathbb{R}^{(m+1)\times d_{i}},\qquad i\in[1:d].$
Step 1: Data preparation: Approximation of the values $g_{0}(x_{i})$ and $G(x_{i})$ at the virtual observation points $x_{i}$, $i\in[1:d]$. The data triplets $(x_{ij},u_{ij},x^{+}_{ij})$, $i\in[1:d_{i}]$, are used to compute an approximation
@@
of the matrix $[g_{0}(x_{i})\mid G(x_{i})]$ for each $x_{i}\in\mathcal{X}$ by solving the linear regression problem
@@
[…]
@@
resembling the approach presented, e.g., in (UmlaPohl18, , Section III.B).
Step 2: Interpolation: The interpolation coefficients are computed analogously to the autonomous case, which leads to the following propagation step of an observable $\psi$:
@@
for all $k,l\in[1:d]$ and $j\in[0:m]$, and $\tilde{g}_{j}$ denotes the $j$-th column of $\tilde{G}$. Then, we directly construct the state-space surrogate of system (22), i.e.,
$x^{+}=f^{\varepsilon}(x,u)=g^{\varepsilon}_{0}(x)+G^{\varepsilon}(x)u,$ (25)
using the approximants (24) of the Koopman operator with $\psi(x)=x_{\ell}$, that is,
$g_{0}^{\varepsilon}(x)_{\ell}:=\sum_{i=1}^{d}(\widehat{K}_{0}(x_{\ell})_{\mathcal{X}})_{i}\Phi_{x_{i}}(x),\qquad(G^{\varepsilon}(x))_{\ell,j}:=\sum_{i=1}^{d}(\widehat{K}_{j}(x_{\ell})_{\mathcal{X}})_{i}\Phi_{x_{i}}(x).$
@@
In the following, we derive a uniform error bound, i.e., Assumption 2, (uniform) Lipschitz continuity as stated in Assumption 3, and a proportional error bound (Assumption 1) for the surrogate (25) and the control-affine dynamics (22) on the bounded set $S=\Omega$. The bounds depend on the fill distance $h_{\mathcal{X}}$ and on the cluster radius $r_{\mathcal{X}}$, and can be made arbitrarily tight by appropriately decreasing these two parameters of the proposed data-driven approximant.
The uniform error bound derived in the next theorem extends Theorem A.1 of the extended arXiv version [^7] See the arXiv preprint https://arxiv.org/abs/2412.02811. of BoldPhil24 to control systems by adapting the arguments used in the proof of (BoldPhil24, , Theorem 4.3). In particular, by choosing cluster radius $r_{\mathcal{X}}$ in the upper bound (26) such that the two terms are balanced, we may ensure a uniform upper bound $\eta:=2C_{1}h_{0}^{k+1/2}$ such that Assumption 2 holds.
@@
Let $k\geq 1$ be the smoothness degree of the Wendland kernel function. There exist constants $C_{1},C_{2},h_{0}>0$ such that the error bound
@@
holds for all $(x,u)\in\Omega\times\mathbb{U}$, where the data satisfies Assumption 5 with fill distance $h_{\mathcal{X}}$ and cluster radius $r_{\mathcal{X}}$ satisfying $h_{\mathcal{X}}\leq h_{0}$ and $r_{\mathcal{X}}<h_{\mathcal{X}}/2$, respectively. Thereby, the error bound depends on the cluster size $r_{\mathcal{X}}$, on the fill distance $h_{\mathcal{X}}$ and on $c$, which is defined by
$c:=\max_{i\in[1:d]}\|U_{i}^{\dagger}\|\cdot\Phi_{n,k}^{1/2}(0)\big{(}\max_{v\in\mathbb{R}^{d}:\|v\|_{\infty}\leq 1}v^{\top}{K}_{\mathcal{X}}^{-1}v\big{)}^{1/2}.$
Proof. In a first step, completely analogous to (BoldPhil24, , Proof of Theorem 4.3), we obtain
@@
for $i\in[1:d]$, where $H:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ is defined by $H(x)=[g_{0}(x)\,|\,G(x)]$ and $\overline{u}:=\max\{\|u\|_{\infty}:u\in\mathbb{U}\}$. Due to uniform continuity of the kernel $\mathsf{k}$ on $\overline{\Omega}\times\overline{\Omega}$, we get the estimate
@@
$\leq$ ${c_{\mathsf{k}}}[L_{g_{0}}+L_{G}\overline{u}]\max\{\sqrt{2d_{i}}\|U_{i}^{\dagger}\|\mid i\in[1:d]\}r_{\mathcal{X}}$
@@
$\leq c_{\mathsf{k}}\tilde{c}\max_{i\in[1:d]}\|U_{i}^{\dagger}\|\|K_{\mathcal{X}}^{-1}\|^{2}r_{\mathcal{X}}$
with $\tilde{c}=[L_{g_{0}}+L_{G}\overline{u}]\cdot\max\{\sqrt{2d_{i}}\mid i\in[1:d]\}$.
Next, we prove the assertion or, to be more precise, we show inequality (26) with $h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})$ instead of $h_{\mathcal{X}}^{k+1/2}$, which will turn out to be beneficial in the proof of Proposition 2. Using (Wend04, , Theorem 11.17) with multiindex $|\alpha|=1$ together with (BoldPhil24, , Theorem 3.7), we get the proportional bound
$|(K_{\mathcal{X}}^{-1}K_{g_{j}(\mathcal{X})}K_{\mathcal{X}}^{-1}\varphi_{\mathcal{X}})^{\top}\textbf{k}_{\mathcal{X}}(x)-\varphi\circ g_{j}(x)|\leq\hat{c}h_{\mathcal{X}}^{k-1/2}\|\varphi\|_{\mathbb{H}}\ {\operatorname{dist}(x,\mathcal{X})}$
with $\hat{c}\geq 0$ for all $\varphi\in\mathbb{H}$ and $\textbf{k}_{\mathcal{X}}=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$. Thus, by the triangle inequality, we obtain a bound on the full error:
$|(\widehat{K}_{j}\varphi_{\mathcal{X}})^{\top}\textbf{k}_{\mathcal{X}}(x)-\varphi\circ g_{j}(x)|\leq\Big{(}\hat{c}h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})+c_{\mathsf{k}}\tilde{c}c\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}\Big{)}\|\varphi\|_{\mathbb{H}}.$
Last, choosing the coordinate functions as observables, we may from now on assume that $\varphi$ is linear. Thus,
$\varphi(x^{+})=\varphi(g_{0}(x)+G(x)u)=\varphi(g_{0}(x))+\sum_{j=1}^{m}\varphi(g_{j}(x))u_{i}$
and hence, we get (26) with $C_{1}=\hat{c}\max_{i\in[1:n]}\|x_{i}\|_{\mathbb{H}}$ and $C_{2}=c_{\mathsf{k}}\tilde{c}\max_{i\in[1:n]}\|x_{i}\|_{\mathbb{H}}$, where $x_{i}$, $i\in[1:n]$ denotes the $i$-th coordinate map. $\square$
As can be seen in Theorem 3, the approximation quality depends on the fill distance $h_{\mathcal{X}}$ of the grid $\mathcal{X}$ of cluster points and on the cluster radius $r_{\mathcal{X}}\geq 0$. The choice of the virtual observation points not only influences the first term of the error bound via the fill distance, but also $\|K_{\mathcal{X}}^{-1}\|$ and $c$ in the second term. Depending on the choice of $\mathcal{X}$, the cluster radius $r_{\mathcal{X}}$ has to be chosen sufficiently small to compensate for $C_{2}c\|K_{\mathcal{X}}^{-1}\|$.
In the following lemma we verify Assumption 3 for the kEDMD surrogate model, i.e. we show that the surrogate model is Lipschitz in the first argument, uniformly in $u$ and $\varepsilon$.
@@
Let $k\geq 1$ and $f\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$ with $\sigma_{n,k}:=\frac{n+1}{2}+k$. Then, there exist constants $h_{0},\bar{L}$ such that, if the fill distance $h_{\mathcal{X}}$ and the cluster radius $r_{\mathcal{X}}$ satisfy $h_{\mathcal{X}}\leq h_{0}$ and $r_{\mathcal{X}}\leq h_{\mathcal{X}}/2$, the surrogate $f^{\varepsilon}$ defined by (25) is Lipschitz continuous in $x$ w.r.t. $u\in\mathbb{U}$ with Lipschitz constant $L_{f^{\varepsilon}}\leq\bar{L}$, i.e.
@@
$\|f^{\varepsilon}(x,u)-f^{\varepsilon}(y,u)\|$$\leq$ $\|x-y\|\|J_{f^{\varepsilon},x}(\xi,u)\pm J_{f,x}(\xi,u)\|$
@@
In the following, let $\hat{g}_{i}$ for $i\in[0:m]$ be the approximation of $g_{i}$ obtained with $r_{\mathcal{X}}=0$, i.e. without errors in the first step of the identification algorithm. To compute a bound for $\|J_{f^{\varepsilon},x}(\xi,u)-J_{f,x}(\xi,u)\|$, we first leverage the control-affine structure and, then, apply (Wend04, , Theorem 11.17) with multiindex $|\alpha|=1$ to obtain
@@
where $C_{J_{g_{i}}}$, $i\in[0:m]$, are constants independent of $h_{\mathcal{X}}$ and $r_{\mathcal{X}}$. Analogously to the proof of (BoldPhil24, , Theorem 4.3), letting $u_{0}=1$, we can estimate for $i\in[0:m]$ and $j\in[1:n]$
$\|[\nabla\hat{g}_{ij}(z)-\nabla g^{\varepsilon}_{ij}(z)]u_{i}\|$
$\hskip 14.22636pt\leq\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\cdot\|K_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(z)\|\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}$
$\hskip 14.22636pt\leq\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\Phi_{n,k}(0)\cdot\max_{v\in\mathds{1}}v^{\top}K_{\mathcal{X}}^{-1}v\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}=:c_{\mathcal{X}i},$ (29)
@@
$\widetilde{C}\!+\!C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}\!\!+m\!\!\!\max\limits_{i=[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}\!\!+n\!\sum_{i=0}^{m}c_{\mathcal{X}i}.$
@@
The continuity in Lemma 1 is uniform in the approximation bound $\varepsilon>0$. This is in contrast to finite-element dictionaries, see schaller2023towards , where the derivative of the ansatz functions increases for decreasing mesh size.
The uniform error estimate in Theorem 3 still contains a constant offset in the upper bound, which results from Step 1 of the approximation scheme. Given an exact approximation of $g_{0}$ and $G$ at the grid points in $\mathcal{X}$, i.e., $r_{\mathcal{X}}=0$, the error bound (26) is proportional in the distance to the grid points. As $x_{1}=0$ due to Assumption 5 the approximation error (26) is bounded by
@@
In the following, we show how one can obtain a proportional error bound even if $r_{\mathcal{X}}>0$ holds. To this end, we leverage the fact that we used the knowledge of the drift dynamics in the origin, that is, we enforced $g_{0}^{\varepsilon}(0)=0$ in the first step of the algorithm, so that the model exactly describes the system in the origin. This is done in the following Proposition.
@@
Let the conditions of Theorem 3 hold. In addition, let $f\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega,\mathbb{R}^{n})$ with $\sigma_{n,k}:=\frac{n+1}{2}+k$ be given. Then, there exist constants $h_{0},r_{0}$ such that, if the fill distance $h_{\mathcal{X}}$ and the cluster radius $r_{\mathcal{X}}$ satisfy $h_{\mathcal{X}}\leq h_{0}$ and $r_{\mathcal{X}}\leq r_{0}$, we have the proportional error bound
@@
where $c_{u}^{\varepsilon}:=C_{1}h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})+C_{2}c\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}$, and
$c_{x}^{\varepsilon}:=C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}+m\max\limits_{i=[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}+n\sum_{i=0}^{m}c_{\mathcal{X}i},$
@@
$\|J_{f^{\varepsilon},x}(x,u)-J_{f,x}(x,u)\|\leq c_{x}^{\varepsilon}.$

### v2

@@ 4.2. Extension of kernel EDMD to control systems
In this section, we recap the extension of kEDMD as introduced in Subsection 4.1 to systems with inputs introduced in Bold et al. (2025b, c). The kernel-EDMD surrogate model is defined for control-affine discrete-time systems
@@
with locally Lipschitz-continuous maps $g_{0}:\Omega\rightarrow\mathbb{R}^{n}$ and $G:\Omega\rightarrow\mathbb{R}^{n\times m}$ with Lipschitz constants $L_{g_{0}},L_{G}$ on $\Omega$ and control input $u$ restricted to a compact set $\mathbb{U}\subset\mathbb{R}^{m}$ containing the origin in its interior. We denote by $g_{i}:\Omega\rightarrow\mathbb{R}^{n}$ the $i$-th column of the matrix-valued map $G$. Moreover, $g_{0}(0)=0$ is assumed to render the origin an equilibrium of the system (19) for $u=0$. For the relation to continuous-time systems, we refer the interested reader to (Bold et al., 2025b, Rem. 4.2) and Strässer et al. (2025), where it is shown that the uniform and proportional error bounds presented below are preserved for fixed $\varepsilon$ and sufficiently small time step $\Delta t>0$ if only the continuous-time system is control affine. Moreover, incorporating minor modifications analogously to Strässer et al. (2026a) leads to a bilinear data-driven surrogate model as shown in Strässer et al. (2025), see also Strässer et al. (2026b).
An alternative data-driven approximation of (19) that shares the advantage of flexible state-control sampling was proposed in Bevanda et al. (2024) using an additional kernel function to express the dependency on the control input. Therein, however, no error bounds were provided. In addition to rigorous and uniform error bounds, which are crucial for provable guarantees for data-driven MPC, the approach proposed in Bold et al. (2025b); Schmitz et al. (2025) allows to counteract numerical ill-conditioning by proposing a two-stage approximation process as shown in the following.
Assumption 3 (Data requirements).
@@
$U_{i}:=\begin{bmatrix}1&\dots&1\\ u_{i1}&\dots&u_{id_{i}}\end{bmatrix}\in\mathbb{R}^{(m+1)\times d_{i}}.$
with $\operatorname{rank}(U_{i})=m+1$ for all $i\in[1:d]$.
Assumption 3 requires $D:=\sum_{i=1}^{d}d_{i}$ data triplets. We call $x_{i}$ a virtual observation (or cluster) point to emphasize that no samples at $x_{i}$ are required. Further, $r_{\mathcal{X}}$ corresponds to the cluster radius, i.e., the radius of the neighborhoods around the virtual observation points, from which the data points are sampled.
[…]
@@
of the matrix $g_{0}(x_{i})\mid G(x_{i})]$ for each $x_{i}\in\mathcal{X}$ by solving the linear regression problem
@@
The solution may be expressed by the pseudoinverse $U_{i}^{\dagger}$, which, for every virtual observation point $x_{i}$, is well defined in view of Assumption 3. Later, we will become interested in a uniform bound on $\max_{i\in[1:d]}\sqrt{d_{i}}\|U_{i}^{\dagger}\|$. On the one hand, this may be achieved by generating (sufficiently) exciting control values $u_{ij}$, $j\in[1:d_{i}]$, following the guidelines derived in Schmitz et al. (2025) leveraging, among others, subspace-angle conditions, to ensure data efficiency ($d_{i}=m+(1-\delta_{1i})$). On the other hand, one may also use randomly chosen input values in combination with a slightly larger $d_{i}$, see (Bold et al., 2025b, Remark 4.6).
@@
resembling the approach presented, e.g., in (Umlauft et al., 2018, Section III.B).
Step 2: Interpolation. The interpolation coefficients are computed analogously to the autonomous case, which leads to the following propagation step of an observable $\psi$:
@@
for all $k,l\in[1:d]$ and all $j\in[0:m]$, where $\tilde{g}_{j}$ denotes the $j$-th column of $\tilde{G}$. Then, we directly construct the state-space surrogate of the control-affine system (19), i.e.,
$x^{+}=f^{\varepsilon}(x,u)\,=\begin{pmatrix}(\psi_{1})_{\varepsilon}^{+}(x,u)\\ \vdots\\ (\psi_{n})_{\varepsilon}^{+}(x,u)\end{pmatrix}=g^{\varepsilon}_{0}(x)+G^{\varepsilon}(x)u,$ (21)
using the approximants (20) of the Koopman operator with observables $\psi_{\ell}(x)=x_{\ell}$ for $\ell\in\{1,\dots,n\}$, that is,
$g_{0}^{\varepsilon}(x)_{\ell}$ $:=\sum_{i=1}^{d}(\widehat{K}_{0}(x_{\ell})_{\mathcal{X}})_{i}\Phi_{x_{i}}(x),$
$(G^{\varepsilon}(x))_{\ell,j}$ $:=\sum_{i=1}^{d}(\widehat{K}_{j}(x_{\ell})_{\mathcal{X}})_{i}\Phi_{x_{i}}(x).$
@@
In the following, we verify Assumptions 1 and 2 for the proposed Koopman surrogate model (21) of the control-affine dynamics (19). To this end, we derive the inequalities (U-bound) and (P-bound) and establish (uniform) Lipschitz continuity on the bounded set $\Omega$. The bounds depend on the fill distance $h_{\mathcal{X}}$ and on the cluster radius $r_{\mathcal{X}}$, and can be made arbitrarily tight by appropriately decreasing these two parameters of the proposed data-driven approximant (21).
The uniform error bound derived in the next theorem for the surrogate (21) extends Theorem A.1 of the extended arXiv version of Bold et al. (2025b) to control systems by adapting the arguments used in the proof of (Bold et al., 2025b, Theorem 4.3). [^7] See https://arxiv.org/abs/2412.02811.
@@
Let $k\geq 1$ be the smoothness degree of the Wendland kernel function. Further, let the dynamics (19) satisfy $g_{0}\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$, $G\in\mathcal{C}_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n\times m})$ with $\sigma_{n,k}:=\frac{n+1}{2}+k$. Suppose that the data satisfies Assumption 3 with fill distance $h_{\mathcal{X}}$ and cluster radius $r_{\mathcal{X}}$ satisfying $h_{\mathcal{X}}\leq h_{0}$ for some $h_{0}>0$ and $r_{\mathcal{X}}<h_{\mathcal{X}}/2$, respectively. Then, there exist constants $C_{1},C_{2}>0$ such that the error bound
@@
holds for all $(x,u)\in\Omega\times\mathbb{U}$, where $c$ is defined by
$c:=\max_{i\in[1:d]}\sqrt{d_{i}}\|U_{i}^{\dagger}\|\cdot\Phi_{n,k}^{1/2}(0)\max_{v\in\mathds{1}}\sqrt{v^{\top}{K}_{\mathcal{X}}^{-1}v}$
and $\mathds{1}=\{v\in\mathbb{R}^{d}\mid v_{i}\in\{\pm 1\}\}$.
Proof. In a first step, analogously to (Bold et al., 2025b, Proof of Theorem 4.2), we obtain
@@
for $i\in[1:d]$, where $H:\mathbb{R}^{n}\to\mathbb{R}^{n\times(m+1)}$ is defined by $H(x)=[g_{0}(x)\,|\,G(x)]$ and $\overline{u}:=\max\{\|u\|_{\infty}:u\in\mathbb{U}\}$. Due to uniform continuity of the kernel $\mathsf{k}$ on $\overline{\Omega}\times\overline{\Omega}$, we get the estimate
@@
$\leq$ ${c_{\mathsf{k}}}(L_{g_{0}}+L_{G}\overline{u})\cdot\max\{\sqrt{2d_{i}}\|U_{i}^{\dagger}\|\mid i\in[1:d]\}r_{\mathcal{X}}$
@@
$\leq\tilde{c}\|K_{\mathcal{X}}^{-1}\|^{2}\max_{i\in[1:d]}\sqrt{d_{i}}\|U_{i}^{\dagger}\|\cdot r_{\mathcal{X}}$
with $\tilde{c}=\sqrt{2}c_{\mathsf{k}}\cdot(L_{g_{0}}+L_{G}\overline{u})$.
Next, we show inequality (22) with $h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})$ instead of $h_{\mathcal{X}}^{k+1/2}$, which will turn out to be beneficial in the proof of Proposition 2. Using (Wendland, 2004, Theorem 11.17) with multiindex $|\alpha|=1$ together with (Bold et al., 2025b, Theorem 3.7), we get the proportional bound
$|(K_{\mathcal{X}}^{-1}K_{g_{j}(\mathcal{X})}K_{\mathcal{X}}^{-1}\varphi_{\mathcal{X}})^{\top}\textbf{k}_{\mathcal{X}}(x)-\varphi\circ g_{j}(x)|$
$\leq$ $\hat{c}h_{\mathcal{X}}^{k-1/2}\|\varphi\|_{\mathbb{H}}\ {\operatorname{dist}(x,\mathcal{X})}$
with $\hat{c}\geq 0$ for all $\varphi\in\mathbb{H}$ and $\textbf{k}_{\mathcal{X}}=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$. Thus, by the triangle inequality, we obtain the following bound on the full error:
$|(\widehat{K}_{j}\varphi_{\mathcal{X}})^{\top}\textbf{k}_{\mathcal{X}}(x)-\varphi\circ g_{j}(x)|$
$\leq$ $\Big(\hat{c}h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})+\tilde{c}c\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}\Big)\|\varphi\|_{\mathbb{H}}.$
Here, the terms $\Phi_{n,k}^{1/2}(0)$ and $\max_{v\in\mathds{1}}\sqrt{v^{\top}{K}_{\mathcal{X}}^{-1}v}$ in $c$ originate from the bound on the norm of $\varphi_{\mathcal{X}}$ and $\mathbf{k}_{\mathcal{X}}(x)$, respectively, similar to (Bold et al., 2025b, Proof of Theorem 4.2). Last, choosing the coordinate functions as observables, we may, for the remainder of the proof, assume that $\varphi$ is linear since we choose $\varphi(x)=x_{i}$ with finite RKHS norm $\|\varphi\|_{\mathbb{H}}$ due to the Sobolev characterization of the Wendland RKHS. Thus,
$\varphi(x^{+})=\varphi(g_{0}(x)+G(x)u)=\varphi(g_{0}(x))+\sum_{j=1}^{m}\varphi(g_{j}(x))u_{j}$
and hence, we get (22) with $C_{1}=\hat{c}\max_{i\in[1:n]}\|x_{i}\|_{\mathbb{H}}$ and $C_{2}=\tilde{c}\max_{i\in[1:n]}\|x_{i}\|_{\mathbb{H}}$, where $x_{i}$, $i\in[1:n]$ denotes the $i$-th coordinate map. $\square$
As can be seen in Theorem 3, the approximation quality mainly depends on the fill distance $h_{\mathcal{X}}$ of the grid $\mathcal{X}$ of cluster points and the cluster radius $r_{\mathcal{X}}\geq 0$. The choice of the virtual observation points $\mathcal{X}$ not only influences the first term of the error bound via the fill distance, but also $\|K_{\mathcal{X}}^{-1}\|$ and $c$ in the second term. Depending on the choice of $\mathcal{X}$, the cluster radius $r_{\mathcal{X}}$ has to be chosen sufficiently small to compensate for $C_{2}c\|K_{\mathcal{X}}^{-1}\|$. Hence, by choosing the cluster radius $r_{\mathcal{X}}$ such that the two terms are balanced, we may ensure the uniform upper bound $\eta^{\varepsilon}:=2C_{1}h_{0}^{k+1/2}$ such that (U-bound) holds.
In the following lemma we verify Assumption 2 for the kEDMD surrogate model, i.e., we show that the surrogate model is Lipschitz in the first argument, uniformly in $u$ and $\varepsilon$.
@@
Let the assumptions of Theorem 3 hold. Then, there exist constants $h_{0},\bar{L}$ such that, if the fill distance $h_{\mathcal{X}}$ and the cluster radius $r_{\mathcal{X}}$ satisfy $h_{\mathcal{X}}\leq h_{0}$ and $r_{\mathcal{X}}\leq h_{\mathcal{X}}/2$, the surrogate $f^{\varepsilon}$ defined by (21) is Lipschitz continuous in $x$ uniformly in $u$ with Lipschitz constant $L_{f^{\varepsilon}}\leq\bar{L}$, i.e.
@@
$\|f^{\varepsilon}(x,u)-f^{\varepsilon}(y,u)\|$
$\leq$ $\|x-y\|\|J_{f^{\varepsilon},x}(\xi,u)\pm J_{f,x}(\xi,u)\|$
@@
In the following, let $\hat{g}_{i}$ for $i\in[0:m]$ be the approximation of $g_{i}$ obtained with $r_{\mathcal{X}}=0$, i.e. without errors in the first step of the identification algorithm. To compute a bound for $\|J_{f^{\varepsilon},x}(\xi,u)-J_{f,x}(\xi,u)\|$, we first leverage the control-affine structure and, then, apply (Wendland, 2004, Theorem 11.17) with multiindex $|\alpha|=1$ to obtain
@@
where $C_{J_{g_{i}}}$, $i\in[0:m]$, are constants independent of $h_{\mathcal{X}}$ and $r_{\mathcal{X}}$. Analogously to the proof of (Bold et al., 2025b, Theorem 4.3), letting $u_{0}=1$, we can estimate for $i\in[0:m]$ and $j\in[1:n]$
$\|[\nabla\hat{g}_{ij}(z)-\nabla g^{\varepsilon}_{ij}(z)]u_{i}\|$ (25)
$\leq$ $\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\cdot\|K_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(z)\|\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}$
$\leq$ $\max_{u\in\mathbb{U}}|u_{i}|L_{J_{g_{i}}}\sqrt{2\max_{k\in[1:d]}d_{k}}\left(\max_{\ell\in[1:d]}\|U^{\dagger}_{\ell}\|\right)\Phi_{n,k}(0)\cdot\max_{v\in\mathds{1}}v^{\top}K_{\mathcal{X}}^{-1}v\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}=:c_{\mathcal{X}i},$
@@
$\widetilde{C}\!+\!C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}\!\!+m\!\!\!\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}\!\!+n\!\sum_{i=0}^{m}c_{\mathcal{X}i}.$
@@
The continuity in Lemma 1 is uniform in the approximation bound $\varepsilon>0$. This is in contrast to finite-element dictionaries, see Schaller et al. (2023), where the derivative of the ansatz functions increases for decreasing mesh size.
The uniform error estimate in Theorem 3 still contains a constant offset in the upper bound, which results from Step 1 of the approximation scheme. Given an exact approximation of $g_{0}$ and $G$ at the grid points in $\mathcal{X}$, i.e., $r_{\mathcal{X}}=0$, and inspecting the last part of the proof of Theorem 3, one may straightforwardly deduce an error bound (22) that is proportional in the distance to the grid points. More precisely, as $x_{1}=0$ due to Assumption 3, the first term in the approximation error (22) may be replaced by
@@
In the following, we show how to ensure a proportional error bound even for $r_{\mathcal{X}}>0$. To this end, we leverage the fact that we used the knowledge of the drift dynamics in the origin, i.e., we enforced $g_{0}^{\varepsilon}(0)=0$ in the first step of the algorithm so that the model exactly describes the system in the origin.
@@
Let the assumptions of Theorem 3 hold. Then, there exist constants $h_{0},r_{0}$ such that, if the fill distance $h_{\mathcal{X}}$ and the cluster radius $r_{\mathcal{X}}$ satisfy $h_{\mathcal{X}}\leq h_{0}$ and $r_{\mathcal{X}}\leq r_{0}$, we have the proportional error bound
@@
for all $x\in\Omega$, $u\in\mathbb{U}$, where $c_{u}^{\varepsilon}:=C_{1}h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})+C_{2}c\|K_{\mathcal{X}}^{-1}\|r_{\mathcal{X}}$, and
$c_{x}^{\varepsilon}:=C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}+m\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}+n\sum_{i=0}^{m}c_{\mathcal{X}i},$
@@
$\quad\|J_{f^{\varepsilon},x}(x,u)-J_{f,x}(x,u)\|$
$\leq\underbrace{C_{J_{g_{0}}}h_{\mathcal{X}}^{k-1/2}\!+\!m\max\limits_{i\in[1:m],u\in\mathbb{U}}C_{J_{g_{i}}}|u_{i}|h_{\mathcal{X}}^{k-1/2}+n\sum_{i=0}^{m}c_{\mathcal{X}i}.}_{=c_{x}^{\varepsilon}}$